OpenStudy (anonymous):

What is the simplified form of (–3x^3y^2)(5xy^-1)? a) 15x^2/y^2 b) –15x^3y^2 c) –15x^4y d) 15x^4y Check my answer I chose C.

OpenStudy (hba):

Well it's easy \[\huge\ x^a \times x^b=x^{a+b}\]

OpenStudy (hba):

Try doin it :)@breeee612

OpenStudy (jiteshmeghwal9):

\(\LARGE{-3.x^3.y^2.5x.y^{-1}}\) \(\LARGE{(-3.5)x^2.x.y^2.y^-1}\) well its easy for u now :)

OpenStudy (anonymous):

When I worked it out I got -15 but I added all the exponents together and that's how I got C...

OpenStudy (hba):

What do you mean by C ?

OpenStudy (anonymous):

Answer choice c

OpenStudy (skullpatrol):

To multiply two powers having the same base you add the exponents.

OpenStudy (hba):

\[\huge\ x^2 \times x^1=\]

OpenStudy (hba):

See my first comment.

OpenStudy (anonymous):

What is the simplified form of –9m–2n5 × 2m–3n–6? a)–18m^5^n b)-18/m^5^n c)-18m^6/n^30 d)–18m6n30 I think it's c or d

OpenStudy (hba):

Close this post and post it in a new thread :)